OpenStudy (artfis):

How do you solve this? I keep getting 0 = 0. :{ 2x - y = 6 -2x + y = -6

OpenStudy (anonymous):

You can't because all terms cancel out!

OpenStudy (tkhunny):

You can and you did! What you have is a Consistent and Dependent System. There are infinitely many solutions.

OpenStudy (artfis):

Wouldn't the answer be 0? Right?

OpenStudy (artfis):

How about 5x - y = 5 -5x + y = 5? I got the same thing too.

OpenStudy (tkhunny):

No, "0" is not an answer. You need an ordered pair. Rewrite one to a familiar form: y = 2x - 6 Just pick an x value. x = 0 Calculate the corresponding y-value. y = -6. There, (0,-6) is one solution. There are infinitely many more.

OpenStudy (artfis):

Oh ok. That clears it up. Thank you.

OpenStudy (tkhunny):

Can you name another solution?